OpenStudy (anonymous):

Solve the given system of equations by substitution. y=x x-y=18 Thank you.

OpenStudy (anonymous):

its y+x=18 sorrry

OpenStudy (anonymous):

Its a plus my mistake.

OpenStudy (anonymous):

Then: y = x Put it in second: x + y = 18 x + x = 18 2x = 18 Divide by 2 both the sides and find x from here.. and y will be same as x..

OpenStudy (anonymous):

Thank you so much as always.

OpenStudy (anonymous):

Welcome dear..